# An operator splitting analysis of Wasserstein–Fisher–Rao gradient flows

### v1

@@ B.1 Moment ODEs and analytic solutions
This representation however is not ideal as it forces the condition $C_{0}^{-1}\neq C_{\pi}^{-1}$ to ensure invertibility of their difference. Note also that $C_{0}=C_{\pi}$ is a valid initial condition, wherein $\frac{dC_{t}}{dt}=0$ for all $t>0$ and (B.4) remains well-posed. We will consider an alternative representation, by first writing (B.17) in terms of $C_{t}$ using the Woodbury matrix identity,
$(A+UWV)^{-1}=A^{-1}-A^{-1}U(W^{-1}+VA^{-1}U)^{-1}VA^{-1},$ (B.18)
with $A=C_{\pi}^{-1},W=(D_{0}+Z_{t})^{-1},\,U=V=e^{-\Gamma t}$ yields
$C_{t}$ $=C_{\pi}-C_{\pi}e^{-\Gamma t}(D_{0}+Z_{t}+e^{-2\Gamma t}C_{\pi})^{-1}e^{-\Gamma t}C_{\pi}$
$=C_{\pi}+(R_{t}^{-1}X_{t}R_{t}^{-1})^{-1}$
where $R_{t}:=C_{\pi}e^{-\Gamma t}$ and $X_{t}:=(C_{\pi}^{-1}-C_{0}^{-1})^{-1}-Z_{t}-e^{-\Gamma t}C_{\pi}e^{-\Gamma t}$. Then let $Z_{t}=Z_{t}^{M}+\Gamma^{-1}(I-e^{-2\Gamma t})$ with $Z_{t}^{M}:=2\int_{0}^{t}e^{-\Gamma s}M_{s}e^{-\Gamma s}ds$. Since $(C_{\pi}^{-1}-C_{0}^{-1})^{-1}=C_{\pi}(C_{0}-C_{\pi})^{-1}C_{0}$, we have
$R_{t}^{-1}X_{t}R_{t}^{-1}$ $=e^{\Gamma t}C_{\pi}^{-1}((C_{\pi}^{-1}-C_{0}^{-1})^{-1}-Z_{t}-e^{-\Gamma t}C_{\pi}e^{-\Gamma t})C_{\pi}^{-1}e^{\Gamma t}$
$=e^{\Gamma t}((C_{0}-C_{\pi})^{-1}C_{0}C_{\pi}^{-1}-C_{\pi}^{-1}(Z_{t}^{M}+\Gamma^{-1}(I-e^{-2\Gamma t}))C_{\pi}^{-1})e^{\Gamma t}-C_{\pi}^{-1}$
$=e^{\Gamma t}((C_{0}-C_{\pi})^{-1}C_{0}C_{\pi}^{-1}-C_{\pi}^{-2}\Gamma^{-1}-C_{\pi}^{-1}Z_{t}^{M}C_{\pi}^{-1})e^{\Gamma t}-C_{\pi}^{-1}+\Gamma^{-1}C_{\pi}^{-2}$ (B.19)
using that $C_{\pi},\,\Gamma$ and $e^{-\Gamma t}$ commute. We can further simplify using Woodbury identities, beginning with
$-C_{\pi}^{-1}+C_{\pi}^{-2}\Gamma^{-1}$ $=-C_{\pi}^{-1}+C_{\pi}^{-2}(C_{\pi}^{-1}+\tfrac{1}{2}I)^{-1}$
$=-C_{\pi}^{-1}+((C_{\pi}^{-1}+\tfrac{1}{2}I)C_{\pi}^{2})^{-1}$
$=-\frac{1}{2}C_{\pi}^{-1}C_{\pi}\left(I+\tfrac{1}{2}C_{\pi}\right)^{-1}$
$=-(2I+C_{\pi})^{-1}$ (B.20)
where in the second last line, we have used the special case of the Woodbury identity
$(I+A)^{-1}-I=-A(I+A)^{-1}$ (B.21)
with $A=\frac{1}{2}C_{\pi}$. Now consider, by similar reasoning,
$(C_{0}-C_{\pi})^{-1}C_{0}C_{\pi}^{-1}-\Gamma^{-1}C_{\pi}^{-2}$ $=(C_{0}-C_{\pi})^{-1}C_{0}C_{\pi}^{-1}-\left(I+\tfrac{1}{2}C_{\pi}\right)^{-1}C_{\pi}^{-1}$
$=\left((C_{0}-C_{\pi})^{-1}C_{0}-I+\tfrac{1}{2}C_{\pi}(I+\tfrac{1}{2}C_{\pi})^{-1}\right)C_{\pi}^{-1}$
$=\left((C_{0}-C_{\pi})^{-1}C_{0}-I\right)C_{\pi}^{-1}+(2I+C_{\pi})^{-1}$
$=(C_{0}-C_{\pi})^{-1}+(2I+C_{\pi})^{-1}$ (B.22)
again using (B.21), $(I+\frac{1}{2}C_{\pi})^{-1}=-I+\frac{1}{2}C_{\pi}(I+\frac{1}{2}C_{\pi})^{-1}$ in the second and third line, and then using (B.18) in the fourth line with $A=-C_{\pi},U=C_{\pi},V=C_{0},W=(C_{\pi})^{-1}$ so that
$(C_{0}-C_{\pi})^{-1}=-C_{\pi}^{-1}-(C_{\pi}-C_{0})^{-1}C_{0}C_{\pi}^{-1}$ (B.23)
Substituting (B.20) and (B.22) into (B.19) yields
$R_{t}^{-1}X_{t}R_{t}^{-1}=e^{\Gamma t}[(C_{0}-C_{\pi})^{-1}+(2I+C_{\pi})^{-1}-C_{\pi}^{-1}Z_{t}^{M}C_{\pi}^{-1}]e^{\Gamma t}-(2I+C_{\pi})^{-1}$
which then yields the expression
$C_{t}=C_{\pi}+(e^{\Gamma t}[(C_{0}-C_{\pi})^{-1}+(2I+C_{\pi})^{-1}-C_{\pi}^{-1}Z_{t}^{M}C_{\pi}^{-1}]e^{\Gamma t}-(2I+C_{\pi})^{-1})^{-1}.$
This form will prove useful in the analysis of $\mathrm{KL}$ decay in Proposition 3.2. Finally, with another application of the Woodbury identity (B.18) with $A=(C_{0}-C_{\pi})^{-1},\,U=V=I,\,W=(2I+C_{\pi})^{-1}-C_{\pi}^{-1}Z_{t}^{M}C_{\pi}^{-1}-e^{-\Gamma t}(2I+C_{\pi})^{-1}e^{-\Gamma t}$, we obtain the representation
$C_{t}$ $=C_{\pi}+(e^{\Gamma t}[(C_{0}-C_{\pi})^{-1}+(2I+C_{\pi})^{-1}-C_{\pi}^{-1}Z_{t}^{M}C_{\pi}^{-1}-e^{-\Gamma t}(2I+C_{\pi})^{-1}e^{-\Gamma t}]e^{\Gamma t})^{-1}$
$=C_{\pi}+e^{-\Gamma t}[(C_{0}-C_{\pi})^{-1}+(2I+C_{\pi})^{-1}-C_{\pi}^{-1}Z_{t}^{M}C_{\pi}^{-1}-e^{-\Gamma t}(2I+C_{\pi})^{-1}e^{-\Gamma t}]^{-1}e^{-\Gamma t}$
$=C_{\pi}+e^{-\Gamma t}((C_{0}-C_{\pi})-(C_{0}-C_{\pi})(W^{-1}+(C_{0}-C_{\pi}))^{-1}(C_{0}-C_{\pi}))e^{-\Gamma t}$
as claimed.
B.2 Derivation of (2.20)
Using the representation (2.14), we will start by calculating the form of $g_{k}$ for $k=2,3,\dots$, and recall from Proposition 2.3 that
$\partial_{\gamma}\eta_{\gamma}=f_{\textrm{W}}(\eta_{\gamma})+f_{\textrm{FR}}(\eta_{\gamma})+\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!}\mathcal{G}_{k}\eta_{\gamma}$
where
$\mathcal{G}_{k}\eta$ $=\eta\left(g_{k}(\eta)-\mathbb{E}_{\eta}[g_{k}]\right)$
$g_{k}(\eta)$ $=-g_{k-1}^{\prime}f_{W}(\eta)+\nabla\log\frac{\eta}{\pi}\cdot\nabla g_{k-1}(\eta)+\frac{1}{\eta}\nabla\cdot(\eta\nabla g_{k-1}(\eta))$
$=-g_{k-1}^{\prime}f_{W}(\eta)+(2\nabla g(\eta)+\nabla\log\pi)^{\top}\nabla g_{k-1}(\eta)+\Delta g_{k-1}(\eta),\enskip k=1,2,3,\dots$
$g_{0}(\eta)$ $=-\log\frac{\eta}{\pi}=-g(\eta)$
where $g$ is as defined in (2.4).
[…]
For $g_{1}$, we have that the Fréchet derivative term is given by $g_{0}^{\prime}f_{\textrm{W}}(\eta)=-\frac{f_{\textrm{W}}}{\eta}$ so that
$g_{1}=\frac{f_{\textrm{W}}(\eta)}{\eta}-|\nabla g_{0}|^{2}-\frac{f_{\textrm{W}}(\eta)}{\eta}=-|\nabla g|^{2}$
For $g_{2}$,
$\nabla g_{1}=-2H\nabla g\,;\quad\Delta g_{1}=-2Tr[H^{2}]$
and the Fréchet derivative term is given by
$g_{1}^{\prime}h$ $=-2\left(\nabla\frac{h}{\eta}\right)^{\top}\nabla g$
$g_{1}^{\prime}f_{\textrm{W}}$ $=-2\left(\nabla\frac{\nabla\cdot(\eta\nabla g)}{\eta}\right)^{\top}\nabla g$
$=-2\left(\nabla(\nabla\log\eta^{\top}\nabla g+\Delta g)\right)^{\top}\nabla g$
$=-2\left(\nabla(\nabla g^{\top}\nabla g+\nabla\log\pi^{\top}\nabla g)\right)^{\top}\nabla g$
$=-2\left(2H\nabla g+H_{\pi}\nabla g+H\nabla\log\pi\right)^{\top}\nabla g$
$=-2\left(-\nabla g_{1}+H_{\pi}\nabla g+H\nabla\log\pi\right)^{\top}\nabla g$
since $\Delta g=Tr[H]$. Combining,
$g_{2}$ $\stackrel{{\scriptstyle+}}{{=}}2\left(-\nabla g_{1}+H_{\pi}\nabla g+H\nabla\log\pi\right)^{\top}\nabla g+(2\nabla g+\nabla\log\pi)^{\top}\nabla g_{1}$
$=2\nabla g^{\top}H_{\pi}\nabla g$
Similarly, for $g_{3}$,
$\nabla g_{2}=4HH_{\pi}\nabla g\,;\quad\Delta g_{2}=4Tr[HH_{\pi}H]$
using the identity $\nabla{\bf v}^{\top}A{\bf w}=\nabla{\bf v}^{\top}A{\bf w}+\nabla{\bf w}^{\top}A^{\top}{\bf v}$ for any ${\bf v}(x),{\bf w}(x):\mathbb{R}^{d\times 1}\rightarrow\mathbb{R}^{d\times 1}$ and $A\in\mathbb{R}^{d\times d}$ constant matrix independent of $x$ (see (93) in [Pedersen2012]). Then for the Fréchet derivative,
$g_{2}^{\prime}h$ $=4\left(\nabla\frac{h}{\eta}\right)^{\top}H_{\pi}\nabla g$
$g_{2}^{\prime}f_{\textrm{W}}$ $=4\left(\nabla\frac{h}{f_{\textrm{W}}}\right)^{\top}H_{\pi}\nabla g=4\left(-\nabla g_{1}+H_{\pi}\nabla g+H\nabla\log\pi\right)^{\top}H_{\pi}\nabla g$
by the same reasoning as earlier. Combining,
$g_{3}$ $\stackrel{{\scriptstyle+}}{{=}}-4\left(-\nabla g_{1}+H_{\pi}\nabla g+H\nabla\log\pi\right)^{\top}H_{\pi}\nabla g+(2\nabla g+\nabla\log\pi)^{\top}\nabla g_{2}$
$=-4\left(2H\nabla g+H_{\pi}\nabla g+H\nabla\log\pi\right)^{\top}H_{\pi}\nabla g+(2\nabla g+\nabla\log\pi)^{\top}4HH_{\pi}\nabla g$
$=\nabla g^{\top}(-8HH_{\pi}-4H_{\pi}^{2}+8HH_{\pi})\nabla g+\nabla\log\pi^{\top}(-4HH_{\pi}+4HH_{\pi})\nabla g$
$=-4\nabla g^{\top}H_{\pi}^{2}\nabla g$
and finally for $g_{4}$, by similar line of reasoning as earlier,
$\nabla g_{3}=-8HH_{\pi}^{2}\nabla g\,;\quad\Delta g_{3}=-8Tr[HH_{\pi}^{2}H]$
and the Fréchet derivative term is given by,
$g_{3}^{\prime}h$ $=-8\left(\nabla\frac{h}{\mu}\right)^{\top}H_{\pi}^{2}\nabla g$
$g_{3}^{\prime}f_{\textrm{W}}$ $=-8\left(-\nabla g_{1}+H_{\pi}\nabla g+H\nabla\log\pi\right)^{\top}H_{\pi}^{2}\nabla g$
Combining,
$g_{4}$ $\stackrel{{\scriptstyle+}}{{=}}8\left(-\nabla g_{1}+H_{\pi}\nabla g+H\nabla\log\pi\right)^{\top}H_{\pi}^{2}\nabla g+(2\nabla g+\nabla\log\pi)^{\top}\nabla g_{3}$
$=8\left(2H\nabla g+H_{\pi}\nabla g+H\nabla\log\pi\right)^{\top}H_{\pi}^{2}\nabla g-8(2\nabla g+\nabla\log\pi)^{\top}HH_{\pi}^{2}\nabla g$
$=8\nabla g^{\top}H_{\pi}^{3}\nabla g$
By induction, it holds that
$g_{k}$ $=-\nabla g^{\top}(-2H_{\pi})^{k-1}\nabla g=-\nabla g^{\top}(2C_{\pi}^{-1})^{k-1}\nabla g$
Finally, we have by the series representation of a matrix exponential that
$\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!}\mathcal{G}_{k}\eta_{k}$ $=-\eta\nabla g^{\top}\left(\frac{C_{\pi}}{2}\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!}(2C_{\pi}^{-1})^{k}\right)\nabla g+\eta\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!}\int\nabla g^{\top}(2C_{\pi}^{-1})^{k-1}\nabla g\eta dx$
$=-\eta\nabla g^{\top}\left(\frac{C_{\pi}}{2}\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!}(2C_{\pi}^{-1})^{k}\right)\nabla g+\eta\int\nabla g^{\top}\left(\frac{C_{\pi}}{2}\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!}(2C_{\pi}^{-1})^{k}\right)\nabla g\eta dx$
$=-\eta\left(g_{\text{FRW}}-\mathbb{E}_{\eta}[g_{\text{FRW}}]\right)$
where
$g_{\text{FRW}}:=\nabla g^{\top}\frac{C_{\pi}}{2}\left(e^{2\gamma C_{\pi}^{-1}}-I\right)\nabla g.$
B.3 Proof of Lemma 3.1
@@
$\mathrm{KL}(\nu_{n}||\pi)$ $=\frac{1}{2}\left[-\log\frac{\det Q_{n\gamma}}{\det C_{\pi}}+(m_{n}^{\beta}-m_{\pi})^{\top}C_{\pi}^{-1}(m_{n}^{\beta}-m_{\pi})+Tr[C_{\pi}^{-1}Q_{n\gamma}]-d\right]$ (B.24)
@@
where $E_{n}^{\beta}$ and $\varepsilon_{n}^{\beta}$ are as defined in (3.7) and (3.8) respectively, $\det(I+C_{\pi}^{-1}E_{n}^{\beta})>0$ and since $\det(A+B)=\det(A)\det(I+A^{-1}B)$ for square matrices $A,B$ with $A$ invertible. A similar expression holds for both $\mathrm{KL}(\mu_{n\gamma}||\pi)$ and $\mathrm{KL}(\eta_{n\gamma}||\pi)$ with $Q_{n\gamma}$ replaced by $C_{n\gamma}$ and $P_{n\gamma}$ respectively.
@@
$Q_{(n+1)\gamma}$ $=C_{\pi}+(e^{\Gamma\gamma}(Q_{n\gamma}-C_{\pi})^{-1}e^{\Gamma\gamma}+(e^{\gamma}-1)C_{\pi}^{-1})^{-1},$ (B.25)
@@
$P_{(n+1)\gamma}=C_{\pi}+\left(e^{\Gamma\gamma}(P_{n\gamma}-C_{\pi})^{-1}e^{\Gamma\gamma}+(e^{\gamma}-1)C_{\pi}^{-1}e^{2\gamma C_{\pi}^{-1}}\right)^{-1}$ (B.26)
Compare these with the recursion obtained from the solving the WFR PDE as derived in (3.1) over time interval $n\gamma,(n+1)\gamma]$, (N.B. in the below, we drop the $\gamma$ from the subscript where there is no ambiguity),
$C_{(n+1)\gamma}=C_{\pi}+(e^{\Gamma\gamma}[(C_{n\gamma}-C_{\pi})^{-1}+(2I+C_{\pi})^{-1}]e^{\Gamma\gamma}-(2I+C_{\pi})^{-1})^{-1}.$ (B.27)
Rearranging (B.25), (B.26) and (B.27) yields
@@
$=\left(e^{n\gamma\Gamma}E_{0}^{-1}e^{n\gamma\Gamma}+(1-e^{\gamma})C_{\pi}^{-1}(I-e^{2\gamma\Gamma})^{-1}(e^{2\gamma n\Gamma}-I)\right)^{-1}$
$=e^{-n\gamma\Gamma}\left(E_{0}^{-1}+(1-e^{\gamma})(I-e^{2\gamma\Gamma})^{-1}C_{\pi}^{-1}(I-e^{-2\gamma n\Gamma})\right)^{-1}e^{-n\gamma\Gamma}$ (B.28)
$E_{n}^{\alpha}$ $=e^{-n\gamma\Gamma}\left(E_{0}^{-1}+(1-e^{\gamma})(e^{-2\gamma C_{\pi}^{-1}}-e^{\gamma})^{-1}C_{\pi}^{-1}(I-e^{-2\gamma n\Gamma})\right)^{-1}e^{-n\gamma\Gamma}$ (B.29)
$E_{n}$ $=\left(e^{n\gamma\Gamma}E_{0}^{-1}e^{n\gamma\Gamma}+\frac{1}{2}\Gamma^{-1}C_{\pi}^{-1}(e^{2n\gamma\Gamma}-I)\right)^{-1}$
$=e^{-n\gamma\Gamma}\left(E_{0}^{-1}+\frac{1}{2}\Gamma^{-1}C_{\pi}^{-1}(I-e^{-2n\gamma\Gamma})\right)^{-1}e^{-n\gamma\Gamma}$ (B.30)
since $\Gamma,e^{\Gamma},C_{\pi}$ all commute. In (B.28) and (B.29) we have used that $\sum_{j=0}^{n-1}T^{j}=(I-T)^{-1}(I-T^{n})$ for (IT)(I-T) invertible with $T=e^{2\gamma\Gamma}$.
@@
$\varepsilon_{n}^{\beta}$ $=E_{n}^{\beta}e^{\gamma C_{\pi}^{-1}}(E_{n-1}^{\beta})^{-1}\varepsilon_{n-1}^{\beta}$
$\varepsilon_{n}^{\alpha}$ $=E_{n}^{\alpha}e^{\gamma C_{\pi}^{-1}}(E_{n-1}^{\alpha})^{-1}\varepsilon_{n-1}^{\alpha}$
$\varepsilon_{n}$ $=E_{n}e^{\gamma C_{\pi}^{-1}}E_{n-1}^{-1}\varepsilon_{n-1}$
@@
and the same expression holds for $(\varepsilon_{n}^{\alpha})^{\top}C_{\pi}^{-1}\varepsilon_{n}^{\alpha}$ and $(\varepsilon_{n})^{\top}C_{\pi}^{-1}\varepsilon_{n}$ with $E_{n}^{\beta}$ replaced by $E_{n}^{\alpha}$ and $E_{n}$ respectively.
Combining all the derived expressions yields the result.
B.4 Proof of Proposition 3.2
@@
(i). Begin with the case where $E_{0}$ is strictly positive definite. Recall $E_{n}=e^{-n\gamma\Gamma}\left(E_{0}^{-1}+\Omega C_{\pi}^{-1}(I-e^{-2n\gamma\Gamma})\right)^{-1}e^{-n\gamma\Gamma}$. It trivially holds that $E_{n}\succ 0$ for all $n=1,2,3\dots$ since $AMA\succ 0$ if $A,M\succ 0$, which is true when letting $A=e^{-n\gamma\Gamma}$ and $M=E_{0}^{-1}+\Omega C_{\pi}^{-1}(I-e^{-2n\gamma\Gamma})$, where the latter is symmetric as $E_{0}^{-1}$ is the inverse of a symmetric matrix and $\Omega,C_{\pi},\Gamma$ are symmetric and they all commute with each other. The same reasoning holds for the claim for $E_{n}^{\beta}$ and $E_{n}^{\alpha}$.
(ii). For the second case, notice that when $E_{0}^{-1}+\Omega C_{\pi}^{-1}\prec 0$, then $E_{0}^{-1}+\Omega C_{\pi}^{-1}(I-e^{-2n\gamma\Gamma})\prec 0$ and $(E_{0}^{-1}+\Omega C_{\pi}^{-1}(I-e^{-2n\gamma\Gamma}))^{-1}\prec 0$ since $\Omega\succ 0$. The claim then holds since $AMA\prec 0$ for all positive definite $A,M$ with $A=e^{-n\gamma\Gamma}$ and $M=(E_{0}^{-1}+\Omega C_{\pi}^{-1}(I-e^{-2n\gamma\Gamma}))^{-1}$. The same reasoning holds for $E_{n}^{\beta},E_{n}^{\alpha}$ with $\Omega$ replaced by $\Omega^{\beta}$ and $\Omega^{\alpha}$ respectively.

### v2

@@ B.1 Moment ODEs and analytic solutions
with $Z_{t}:=2\int_{0}^{t}e^{-\Gamma s}(I+M_{s})e^{-\Gamma s}ds$ and since $D_{0}=(C_{0}^{-1}-C_{\pi}^{-1})^{-1}$. Inserting into (2.14) yields an analytic solution for $C_{t}^{-1}$. Note that for the special case $M_{t}=0$, the above coincides with the expression for $C_{t}^{-1}$ in [LMT+25, Remark 2.3] with $\alpha=\beta=1$ there. Note also that $Z_{t}$ is well defined for all $t>0$ for the choices of $M_{t}$ considered in this manuscript, since it is easily checked that the integrand approaches zero asymptotically in $t$. This representation however is not ideal as it forces the condition $C_{0}^{-1}\neq C_{\pi}^{-1}$ to ensure invertibility of their difference. Note also that $C_{0}=C_{\pi}$ is a valid initial condition, wherein $\frac{dC_{t}}{dt}=0$ for all $t>0$ and (2.4) remains well-posed. We will consider an alternative representation, by using Woodbury matrix identity to obtain
$C_{t}$ $=C_{\pi}-C_{\pi}e^{-\Gamma t}(D_{0}+Z_{t}+e^{-2\Gamma t}C_{\pi})^{-1}e^{-\Gamma t}C_{\pi}=:C_{\pi}+(R_{t}^{-1}X_{t}R_{t}^{-1})^{-1}$ (2.15)
where $R_{t}:=C_{\pi}e^{-\Gamma t}$ and $X_{t}:=(C_{\pi}^{-1}-C_{0}^{-1})^{-1}-Z_{t}-e^{-\Gamma t}C_{\pi}e^{-\Gamma t}$. Then let $Z_{t}=Z_{t}^{M}+\Gamma^{-1}(I-e^{-2\Gamma t})$ with $Z_{t}^{M}:=2\int_{0}^{t}e^{-\Gamma s}M_{s}e^{-\Gamma s}ds$. A repeated application of Woodbury matrix identity on $R_{t}^{-1}X_{t}R_{t}^{-1}$ and using that $C_{\pi},\,\Gamma$ and $e^{-\Gamma t}$ commute yields
$R_{t}^{-1}X_{t}R_{t}^{-1}=e^{\Gamma t}[(C_{0}-C_{\pi})^{-1}+(2I+C_{\pi})^{-1}-C_{\pi}^{-1}Z_{t}^{M}C_{\pi}^{-1}]e^{\Gamma t}-(2I+C_{\pi})^{-1}.$
Substituting back into (2.15) yields the result.
B.2 Proof of Lemma 3.1
@@
$\mathrm{KL}(\nu_{n}||\pi)$ $=\frac{1}{2}\left[-\log\frac{\det Q_{n}}{\det C_{\pi}}+(m_{n}^{\beta}-m_{\pi})^{\top}C_{\pi}^{-1}(m_{n}^{\beta}-m_{\pi})+Tr[C_{\pi}^{-1}Q_{n}]-d\right]$ (2.16)
@@
where $E_{n}^{\beta}$ and $\varepsilon_{n}^{\beta}$ are as defined in (3.7) and (3.8) respectively, $\det(I+C_{\pi}^{-1}E_{n}^{\beta})>0$ since $\det(A+B)=\det(A)\det(I+A^{-1}B)$ for square matrices $A,B$ with $A$ invertible. A similar expression holds for both $\mathrm{KL}(\mu_{n\gamma}||\pi)$ and $\mathrm{KL}(\eta_{n}||\pi)$ with $Q_{n}$ replaced by $C_{n\gamma}$ and $P_{n}$ respectively.
@@
$Q_{n+1}$ $=C_{\pi}+(e^{\Gamma\gamma}(Q_{n}-C_{\pi})^{-1}e^{\Gamma\gamma}+(e^{\gamma}-1)C_{\pi}^{-1})^{-1},$ (2.17)
@@
$P_{n+1}=C_{\pi}+\left(e^{\Gamma\gamma}(P_{n}-C_{\pi})^{-1}e^{\Gamma\gamma}+(e^{\gamma}-1)C_{\pi}^{-1}e^{2\gamma C_{\pi}^{-1}}\right)^{-1}$ (2.18)
Compare these with the recursion obtained from the solving the WFR PDE as derived in (3.1) over time interval $n\gamma,(n+1)\gamma]$,
$C_{(n+1)\gamma}=C_{\pi}+(e^{\Gamma\gamma}[(C_{n\gamma}-C_{\pi})^{-1}+(2I+C_{\pi})^{-1}]e^{\Gamma\gamma}-(2I+C_{\pi})^{-1})^{-1}.$ (2.19)
Rearranging (2.17), (2.18) and (2.19) yields
@@
$=e^{-n\gamma\Gamma}\left(E_{0}^{-1}+(1-e^{\gamma})(I-e^{2\gamma\Gamma})^{-1}C_{\pi}^{-1}(I-e^{-2\gamma n\Gamma})\right)^{-1}e^{-n\gamma\Gamma}$ (2.20)
$E_{n}^{\alpha}$ $=e^{-n\gamma\Gamma}\left(E_{0}^{-1}+(1-e^{\gamma})(e^{-2\gamma C_{\pi}^{-1}}-e^{\gamma})^{-1}C_{\pi}^{-1}(I-e^{-2\gamma n\Gamma})\right)^{-1}e^{-n\gamma\Gamma}$ (2.21)
$E_{n}$ $=e^{-n\gamma\Gamma}\left(E_{0}^{-1}+\frac{1}{2}\Gamma^{-1}C_{\pi}^{-1}(I-e^{-2n\gamma\Gamma})\right)^{-1}e^{-n\gamma\Gamma}$ (2.22)
since $\Gamma,e^{\Gamma},C_{\pi}$ all commute. In (2.20) and (2.21) we have used that $\sum_{j=0}^{n-1}T^{j}=(I-T)^{-1}(I-T^{n})$ for (IT)(I-T) invertible with $T=e^{2\gamma\Gamma}$.
@@
$\varepsilon_{n}^{\beta}$ $=E_{n}^{\beta}e^{\gamma C_{\pi}^{-1}}(E_{n-1}^{\beta})^{-1}\varepsilon_{n-1}^{\beta};\qquad\varepsilon_{n}^{\alpha}=E_{n}^{\alpha}e^{\gamma C_{\pi}^{-1}}(E_{n-1}^{\alpha})^{-1}\varepsilon_{n-1}^{\alpha};\qquad\varepsilon_{n}=E_{n}e^{\gamma C_{\pi}^{-1}}E_{n-1}^{-1}\varepsilon_{n-1}$
@@
and the same expression holds for $(\varepsilon_{n}^{\alpha})^{\top}C_{\pi}^{-1}\varepsilon_{n}^{\alpha}$ and $(\varepsilon_{n})^{\top}C_{\pi}^{-1}\varepsilon_{n}$ with $E_{n}^{\beta}$ replaced by $E_{n}^{\alpha}$ and $E_{n}$ respectively. Combining all the derived expressions yields the result.
B.3 Proof of Proposition 3.2
@@
(i) Here we assume $E_{0}$ is strictly positive definite and recall $E_{n}=e^{-n\gamma\Gamma}(E_{0}^{-1}+\Omega V_{n})^{-1}e^{-n\gamma\Gamma}$ where $V_{n}:=C_{\pi}^{-1}(I-e^{-2n\gamma\Gamma})$. It trivially holds that $E_{n}\succ 0$ for all $n=1,2,3\dots$, recalling that $XYX$ is p.d. for any $X,Y$ symmetric p.d. It is clear that $e^{-n\gamma\Gamma}$ and $U_{n}:=E_{0}^{-1}+\Omega V_{n}$ are both symmetric p.d since $E_{0}^{-1}$ is the inverse of a symmetric matrix and $\Omega,C_{\pi},\Gamma$ are symmetric p.d and they all commute with each other. The same reasoning holds for the claim for $E_{n}^{\beta}$ and $E_{n}^{\alpha}$.
For the second part of the claim, first note that $\Omega,\Omega^{\beta},\Omega^{\alpha}$ are p.d. and
$\Omega^{\alpha}\succ\Omega\succ\Omega^{\beta},$ (2.23)
since $\Omega^{-1}\prec(\Omega^{\beta})^{-1}\Leftrightarrow 2C_{\pi}^{-1}\prec\tfrac{e^{\gamma}}{e^{\gamma}-1}(e^{2\gamma C_{\pi}^{-1}}-I),$ which is true for any symmetric p.d. $C_{\pi}$, using the standard matrix exponential inequality, $e^{X}\succ I+X$ for any symmetric p.d. $X$, and that $\tfrac{1}{\gamma}<\frac{1}{1-e^{-\gamma}}$ for all $\gamma>0$. For the first inequality in (2.23), we have
$\Omega^{-1}\succ(\Omega^{\alpha})^{-1}$ $\Leftrightarrow 2C_{\pi}^{-1}+I\succ\tfrac{1}{1-e^{\gamma}}(I-e^{2\gamma\Gamma})e^{-2\gamma C_{\pi}^{-1}}$
$\Leftrightarrow 2(e^{\gamma}-1)C_{\pi}^{-1}\succ I-e^{-2\gamma C_{\pi}^{-1}},$
which is true for any symmetric p.d. $C_{\pi}^{-1}$ and $\gamma>0$ since
$I-e^{-2\gamma C_{\pi}^{-1}}=2\gamma C_{\pi}^{-1}\int_{0}^{1}e^{-2\gamma tC_{\pi}^{-1}}dt\prec 2\gamma C_{\pi}^{-1}\prec 2(e^{\gamma}-1)C_{\pi}^{-1}.$
Then since $\Omega^{\alpha},\Omega,V_{n}$ commute and are all p.d., $(E_{0}^{-1}+\Omega^{\alpha}V_{n})^{-1}\prec(E_{0}^{-1}+\Omega V_{n})^{-1}$. Also, since $e^{-n\gamma\Gamma}\succ 0$, it holds that
$e^{-n\gamma\Gamma}(E_{0}^{-1}+\Omega^{\alpha}V_{n})^{-1}e^{-n\gamma\Gamma}\prec e^{-n\gamma\Gamma}(E_{0}^{-1}+\Omega V_{n})^{-1}e^{-n\gamma\Gamma}$
[…]
(ii) For the second case, first notice that $E_{0}\prec 0$ implies $E_{0}^{-1}+\Omega C_{\pi}^{-1}\prec 0$ since
$(C_{0}-C_{\pi})^{-1}+\Omega C_{\pi}^{-1}\prec 0\Leftrightarrow C_{\pi}-C_{0}\prec C_{\pi}\Omega^{-1}\Leftrightarrow C_{0}\succ-2I$
which holds for any p.d $C_{0}$. Therefore, $MU_{n}\prec 0$ where $U_{n}$ is as defined in (i). Also, for the W-FR split case, $E_{0}\prec 0$ implies $U_{n}^{\beta}:=E_{0}^{-1}+\Omega^{\beta}V_{n}\prec 0$ since
$(C_{0}-C_{\pi})^{-1}+\Omega^{\beta}C_{\pi}^{-1}(I-e^{-2n\gamma\Gamma})\prec 0$ $\Leftrightarrow C_{\pi}-C_{0}\prec(I-e^{-2n\gamma\Gamma})^{-1}C_{\pi}(\Omega^{\beta})^{-1}$
$\Leftrightarrow C_{0}\succ C_{\pi}\left(I-\tfrac{1}{1-e^{\gamma}}(I-e^{-2n\gamma\Gamma})^{-1}(I-e^{2\gamma\Gamma})\right)$
which is true for any p.d $C_{0}$ since $\Gamma\succ\tfrac{1}{2}I$ implies $\tfrac{1}{e^{\gamma}-1}(e^{2\gamma\Gamma}-I)\succ I-e^{-2n\gamma\Gamma}$ from which it follows that $I-\tfrac{1}{1-e^{\gamma}}(I-e^{-2n\gamma\Gamma})^{-1}(I-e^{2\gamma\Gamma})\prec 0$. Similarly, for the FR-W Split scheme, $E_{0}\prec 0$ implies $U_{n}^{\alpha}:=E_{0}^{-1}+\Omega^{\alpha}V_{n}\prec 0$ by the same reasoning as for W-FR. Then by the same reasoning as in (i), $E_{n}\prec 0$ and the same holds for $E_{n}^{\beta}$ and $E_{n}^{\alpha}$.
The second part of the claim follows by combining the above and (2.23) to obtain $0\succ U_{n}^{\alpha}\succ U_{n}\succ U_{n}^{\beta},$ from which it follows that $E_{n}^{\alpha}\succ-E_{n}\succ-E_{n}^{\beta}\succ 0,$ and then using the same arguments for the proof of the second claim of (i).